Prepare the Income Statement of Waltons Manufacturers for the month ended 28 February 2021 using the absorption costing method.

INFORMATION

The following forecasts were obtained from the accounting records of Waltons Manufacturers for the month ended 28 February 2021:

Inventory on 01 February 2021

Nil

Production

25 000 units

Sales

23 000 units

Selling price per unit

R100

Manufacturing costs:

Fixed manufacturing costs

R240 000 per month

Variable manufacturing costs per unit

R42

Marketing costs:

Sales personnel’s salaries and advertising

R40 000 per month

Delivery costs

R5 per unit sold

Administration costs:

Salaries

R50 000 per month

Other office costs

R4 per unit sold


1
Expert's answer
2020-11-03T16:03:01-0500

Sales: "23,000*100=2,300,000"

Cost:

Fixed manufacturing costs: 240,000

Variable manufacturing costs: "42*25,000=1,050,000"

Sales personnel’s salaries and advertising: 40,000

Delivery costs: "5*23,000=115,000"

Salaries: 50,000

Other office costs: "4*23,000=92,000"

Total cost: 1,587,000

Income before taxes: 713,000